OpenStudy (anonymous):

Find $\left[\left(\frac{2}{5}\right)^{12} \left(\frac{3}{16}\right)^{-8}\right] \cdot \left[\left(\frac{2}{5}\right)^{12} \left(\frac{3}{16}\right)^{-8}\right]^{-1}.$

OpenStudy (camerondoherty):

\(\left[\left(\frac{2}{5}\right)^{12} \left(\frac{3}{16}\right)^{-8}\right] \cdot \left[\left(\frac{2}{5}\right)^{12} \left(\frac{3}{16}\right)^{-8}\right]^{-1}.$\)

OpenStudy (camerondoherty):

You mean tht?

OpenStudy (anonymous):

yep